# Splitting obstructions and $\mathbb{Z}_{2}$ invariants in time-reversal symmetric topological insulators

### v1

@@ 1. Setting of the problem: from periodic Hamiltonians to projection-valued maps
$T(k+\lambda_{j}^{*})=\tau_{\lambda_{j}^{*}}T(k)\tau_{\lambda_{j}^{*}}^{-1}.$
If the fibers are constant, this reduces to $Te^{iL_{j}}=e^{-iL_{j}}T$, and by functional calculus one can choose $L_{j}$ such that $TL_{j}=L_{j}T$, so that
@@
When these properties hold simultaneously, we can define a unitary-valued map $U:\mathbb{R}^{d}\to\mathcal{U}(\mathcal{H})$ as
$U(k_{1}^{*},\dots,k_{d}^{*})=\prod_{j=1}^{d}e^{iL_{j}k_{j}^{*}}.$
@@
$TU(k_{1}^{*},\dots,k_{d}^{*})=U(-k_{1}^{*},\dots,-k_{d}^{*})T.$
If $(k_{1}^{*},\dots,k_{d}^{*})=k^{*}$ are the coordinates of $k\in\mathbb{R}^{d}$ in the chosen lattice basis, we define
$\tilde{P}_{\Omega}(k)=U(k_{1}^{*},\dots,k_{d}^{*})P_{\Omega}(k)U(k_{1}^{*},\dots,k_{d}^{*})^{-1}.$
@@
The choice of $L_{1},\dots,L_{d}$ is not unique: if another unitary-valued map $U^{\prime}:\mathbb{R}^{d}\to\mathcal{U}(\mathcal{H})$ has the same properties, then the corresponding projections $\tilde{P}_{\Omega}$ and $\hat{P}_{\Omega}$ are unitarily equivalent via the periodic unitary-valued map $U^{\prime}(k^{*})U(k^{*})^{-1}$:
$\hat{P}_{\Omega}(k)=U^{\prime}(k^{*})U(k^{*})^{-1}\tilde{P}_{\Omega}(k)\left(U^{\prime}(k^{*})U(k^{*})^{-1}\right)^{-1},$
with the time-reversal constraint
$TU^{\prime}(k^{*})U(k^{*})^{-1}=U^{\prime}(-k^{*})U(-k^{*})^{-1}T,\quad\forall k^{*}\in\mathbb{R}^{d}.$
With this construction, we obtain a regular projection-valued map $P:\mathbb{R}^{d}\to\operatorname{Proj}_{m}(\mathcal{H})$ that is $\Lambda^{*}$-periodic in $k$. We may reinterpret it as a projection-valued map $P:\mathbb{T}^{d}\to\operatorname{Proj}_{m}(\mathcal{H})$ since $\mathbb{R}^{d}/\Lambda^{*}\simeq\mathbb{T}^{d}$, preserving the symmetry constraint
@@
Before giving a formal definition, we briefly consider again the discrete setting, where the Bloch-Floquet transform and the procedure to eliminate the $\tau$-covariance can be summarized in a single step. As we will see, we again obtain a projection-valued map that is periodic and satisfies a symmetry constraint analogous to the one above.
Let $X\subset\mathbb{R}^{d}$ be a discrete configuration space with a lattice $\Lambda\simeq\mathbb{Z}^{d}$ such that $X=X+\lambda$ for all $\lambda\in\Lambda$, and suppose $X/\Lambda$ is finite. We can then select a finite set of points $\mathbb{W}=\{x_{1},\dots,x_{l}\}\subset X$ such that $X=\mathbb{W}+\Lambda$. Let $H$ be an Hamiltonian acting in $L^{2}(X)\otimes\mathbb{C}^{n}$, which we rewrite as
$L^{2}(X)\otimes\mathbb{C}^{n}\simeq L^{2}(\mathbb{W}\times\Lambda)\otimes\mathbb{C}^{n}\simeq L^{2}(\Lambda)\otimes\mathbb{C}^{l\cdot n}\simeq L^{2}(\mathbb{Z}^{d})\otimes\mathbb{C}^{l\cdot n}.$
@@
$\mathcal{F}:L^{2}(\mathbb{Z}^{d})\to L^{2}(\mathbb{T}^{d},dk,\mathbb{C}),\quad(\mathcal{F}f)(k)=\sum_{n\in\mathbb{Z}^{d}}\frac{1}{(2\pi)^{d/2}}e^{ik\cdot n}f(n),$
@@
is decomposable. That is, there exists a family of self-adjoint matrices $H(k)$ on $\mathcal{H}=\mathbb{C}^{l\cdot n}$ such that
$\big[(\mathcal{F}\otimes\mathds{1})\mathcal{W}H\mathcal{W}^{-1}(\mathcal{F}\otimes\mathds{1})^{-1}(f)\big]_{k}=H(k)f_{k},\quad\forall f\in\mathcal{H},\ k\in\mathbb{T}^{d}.$
Here, periodicity in $k$ holds directly, which is why we write $k\in\mathbb{T}^{d}$. The same steps as in the continuum case yield a projection-valued map $P:\mathbb{T}^{d}\to\operatorname{Proj}_{m}(\mathcal{H})$ satisfying
@@
It is important to note that the choice of discrete Wigner-Seitz cell $\mathbb{W}=\{x_{1},\dots,x_{n}\}$ is not unique, and the ordering of points could in principle affect the topological properties of the model. If a different cell $\mathbb{W}^{\prime}=\{x_{1}^{\prime},\dots,x_{n}^{\prime}\}$ is chosen, one defines another isomorphism $\mathcal{W}^{\prime}$ and obtains a new fiber decomposition
$\big[(\mathcal{F}\otimes\mathds{1})\mathcal{W}^{\prime}H(\mathcal{W}^{\prime})^{-1}(\mathcal{F}\otimes\mathds{1})^{-1}(f)\big]_{k}=H^{\prime}(k)f_{k},\quad\forall f\in\mathcal{H},\ k\in\mathbb{T}^{d}.$
@@
$(\mathcal{F}\otimes\mathds{1})\mathcal{W}^{\prime}\mathcal{W}^{-1}(\mathcal{F}\otimes\mathds{1})^{-1}=\int_{k\in\mathbb{T}^{d}}^{\oplus}U(k),$
@@
Let $\mathcal{H}$ be a Hilbert space, $T$ an anti-unitary operator with $T^{2}=-\mathds{1}$, $m$ the rank of the projections, and $d$ the physical dimension of the model. A continuous projection-valued map $P:\mathbb{T}^{d}\to\operatorname{Proj}_{m}(\mathcal{H})$ is time-reversal symmetric if
@@
Two such maps $P_{0},P_{1}$ are unitarily equivalent if there exists a continuous unitary-valued map $U:\mathbb{T}^{d}\to\mathcal{U}(\mathcal{H})$ such that
(5) $TU(k)=U(-k)T,\quad P_{1}(k)=U(k)P_{0}(k)U(k)^{-1},\quad\forall k\in\mathbb{T}^{d}.$
@@
Two maps $P_{0},P_{1}$ are homotopic if there exists a continuous projection-valued map $P:[0,1]\times\mathbb{T}^{d}\to\operatorname{Proj}_{m}(\mathcal{H})$ such that
(6) $P(0,k)=P_{0}(k),\quad P(1,k)=P_{1}(k),\quad TP(t,k)=P(t,-k)T,\quad\forall t\in[0,1],\ k\in\mathbb{T}^{d}.$
Remark 1.5 (Continuous vs. Analytic).
Since the family of projectors $\{P(k)\}_{k\in\mathbb{R}^{d}}$ appearing in physical models is usually analytic in $k$, one may ask whether real-analytic Bloch frames can be constructed. Following the argument of Lemma 2.3 in [cornean2016construction], one can show that whenever a continuous Bloch frame exists, it can be modified to a real-analytic one while preserving the time-reversal symmetry. Hence, in the following, we will freely interchange continuous, differentiable, and analytic structures.
Remark 1.6 (Role of Unitary Equivalences and Homotopies).
@@
It is important to emphasize that unitary equivalence is associated with the choice of dimerization in tight-binding models, or with the selection of a unitary-valued map that removes the $\tau$-covariance in continuum models. In contrast, homotopies correspond to continuous deformations of physical parameters in the original Hamiltonian that do not close the spectral gaps. Under such deformations, the topological properties of $P$ are preserved. Since the physical interpretations of these two mathematical notions differ, it is natural to treat them separately. Moreover, when the two conditions coincide mathematically, they establish an equivalence between distinct physical aspects of the system.
2. Preliminary results
In this section, we introduce two constructions that will play a central role throughout the text. The first serves as a substitute for parallel transport in the continuous setting and will be crucial for defining the invariant. The second provides an equivalent formulation of the Chern number, which we will later use to reinterpret the invariant through a suitable decomposition of the projector .
2.1. Continuous parallel transport

### v2

@@ 1. From (symmetric) periodic Hamiltonians to (symmetric) projection-valued maps
$T(k+\lambda_{j}^{*})=\tau_{\lambda_{j}^{*}}T(k)\tau_{\lambda_{j}^{*}}.$
If the fibers are constant, this reduces to $Te^{iL_{j}}=e^{-iL_{j}}T$; by functional calculus one can choose $L_{j}$ such that $TL_{j}=-L_{j}T$, so that
@@
We can now define a unitary-valued map $U:\mathbb{R}^{d}\to\mathcal{U}(\mathcal{H}_{\mathrm{f}})$ as
$U(k_{1},\dots,k_{d})=\prod_{j=1}^{d}e^{iL_{j}k_{j}}.$
@@
$TU(k_{1},\dots,k_{d})=U(-k_{1},\dots,-k_{d})T.$
If $(k_{1},\dots,k_{d})=k$ are the coordinates of $k\in\mathbb{R}^{d}$ in the chosen lattice basis, we define
$\tilde{P}_{\Omega}(k)=U(k_{1},\dots,k_{d})P_{\Omega}(k)U(k_{1},\dots,k_{d})^{-1}\equiv U(k)P_{\Omega}(k)U(k)^{-1}.$
@@
The choice of $L_{1},\dots,L_{d}$ is not unique: if another unitary-valued map $U^{\prime}:\mathbb{R}^{d}\to\mathcal{U}(\mathcal{H}_{\mathrm{f}})$ has the same properties as the $U$ constructed above, then the corresponding projections $\tilde{P}_{\Omega}$ and $\tilde{P}^{\prime}_{\Omega}$ are unitarily equivalent via the periodic unitary-valued map $U^{\prime}(k)U(k)^{-1}$:
$\tilde{P}^{\prime}_{\Omega}(k)=U^{\prime}(k)U(k)^{-1}\tilde{P}_{\Omega}(k)\left[U^{\prime}(k)U(k)^{-1}\right]^{-1},$
with the conjugation map satisfying the time-reversal symmetry constraint
$TU^{\prime}(k)U(k)^{-1}=U^{\prime}(-k)U(-k)^{-1}T,\quad\forall k\in\mathbb{R}^{d}.$
With this construction, we obtain a regular projection-valued map $P:\mathbb{R}^{d}\to\operatorname{Proj}_{m}(\mathcal{H}_{\mathrm{f}})$ that is $\Lambda^{*}$-periodic in $k$. We may reinterpret it as a projection-valued map $P:\mathbb{T}^{d}\to\operatorname{Proj}_{m}(\mathcal{H}_{\mathrm{f}})$ since $\mathbb{R}^{d}/\Lambda^{*}\simeq\mathbb{T}^{d}$, which furthermore satisfies the symmetry constraint
@@
Before giving a formal definition, we briefly consider again the discrete setting, where the Bloch-Floquet transform and the procedure to reduce $\tau$-covariance to $\Lambda^{*}$-periodicity can be summarized in a single step. As we will see, we obtain again a projection-valued map that is periodic and satisfies a symmetry constraint analogous to the one above.
Let $X\subset\mathbb{R}^{d}$ be a discrete configuration space which is invariant under shifts in a lattice $\Lambda\simeq\mathbb{Z}^{d}$, and suppose $X/\Lambda$ is finite. We can then select a finite set of points $\mathbb{W}=\{x_{1},\dots,x_{l}\}\subset X$ such that $X=\mathbb{W}+\Lambda$. Let $H$ be an Hamiltonian acting in $L^{2}(X)\otimes\mathbb{C}^{n}$, which we rewrite as
$L^{2}(X)\otimes\mathbb{C}^{N}\simeq L^{2}(\mathbb{W}\times\Lambda)\otimes\mathbb{C}^{N}\simeq\ell^{2}(\Lambda)\otimes\mathbb{C}^{l\cdot N}\simeq\ell^{2}(\mathbb{Z}^{d})\otimes\mathbb{C}^{l\cdot n}.$
@@
$\mathcal{F}:\ell^{2}(\mathbb{Z}^{d})\to L^{2}(\mathbb{T}^{d},dk,\mathbb{C}),\quad(\mathcal{F}f)(k)=\sum_{\lambda\in\mathbb{Z}^{d}}\frac{1}{(2\pi)^{d/2}}e^{ik\cdot\lambda}f(\lambda),$
@@
is decomposable into a family of self-adjoint fibers $\{H(k)\}_{k\in\mathbb{T}^{d}}$, in the sense specified above. Here, periodicity in $k$ holds directly, which is why we write $k\in\mathbb{T}^{d}$. The same steps as in the continuum case yield a projection-valued map $P:\mathbb{T}^{d}\to\operatorname{Proj}_{m}(\mathbb{C}^{l\cdot n})$ satisfying
@@
It is important to note that the choice of discrete Wigner-Seitz cell $\mathbb{W}=\{x_{1},\dots,x_{l}\}$ is not unique, and this choice could in principle affect the topological properties of the model. If a different cell $\mathbb{W}^{\prime}=\{x_{1}^{\prime},\dots,x_{l}^{\prime}\}$ is chosen, one defines another isomorphism $\mathcal{W}^{\prime}$ and obtains a new fiber decomposition
$\big[(\mathcal{F}\otimes\mathds{1})\mathcal{W}^{\prime}H(\mathcal{W}^{\prime})^{-1}(\mathcal{F}\otimes\mathds{1})^{-1}(f)\big]_{k}=H^{\prime}(k)f_{k},\quad\forall f\in L^{2}(\mathbb{T}^{d},dk,\mathbb{C}^{l\cdot N}),\ k\in\mathbb{T}^{d}.$
@@
$(\mathcal{F}\otimes\mathds{1})\mathcal{W}^{\prime}\mathcal{W}^{-1}(\mathcal{F}\otimes\mathds{1})^{-1}=\int_{\mathbb{T}^{d}}^{\oplus}dk\,U(k),$
@@
Let $d\in\mathbb{N}$ (the physical dimension of the model), $m\in\mathbb{N}$ (the rank of the projections), $\mathcal{H}$ a Hilbert space, and $T$ an anti-unitary operator on $\mathcal{H}$ with $T^{2}=-\mathds{1}$, $m\in\mathbb{N}$. A continuous, periodic projection-valued map $P:\mathbb{T}^{d}\to\operatorname{Proj}_{m}(\mathcal{H})$ is time-reversal symmetric if
@@
Notice in particular that, on points $k_{\star}$ fixed by the involution $k\mapsto-k$ on the torus $\mathbb{T}^{d}$, the antiunitary operator $T$ induces the structure of an $m$-dimensional quaternionic Hilbert space on the range of the projection $P(k_{\star})$, as $TP(k_{\star})=P(k_{\star})T$. Applying Lemma A.1 from Appendix A yields the result that the dimension of $\operatorname{Ran}(P(k_{\star}))$ must be even. Since the rank is constant in $k$, we have that $m\in 2\mathbb{Z}$. So from now on we impose $m=2n$.
Definition 1.5 (Symmetric periodic (Bloch) frame).
A frame for a projection-valued map $\{P(k)\}_{k\in\mathbb{R}^{d}}$ is a collection of orthonormal vectors $\{v_{j}(k)\}_{j\in\{1,\cdots,m\}}$ that are continuous in $k$ and span $\operatorname{Ran}(P(k))$ orthonormally for all $k\in\mathbb{R}^{d}$.
If the projection-valued map $\{P(k)\}_{k\in\mathbb{T}^{d}}$ is $\Lambda$-periodic, the frame is called periodic if the vectors $v_{j}(k)$ depend $\Lambda$-periodically on $k$.
If the projection-valued map is time-reversal symmetric, and therefore $m=2n$, the frame is called symmetric if
$Tv_{j}(k)=v_{j+n}(-k)\quad\text{for all }1\leq j\leq n.$
Notice how the symmetry constraint couples vectors in a frame into Kramers pairs, and it is not possible to relate a vector to itself by imposing Tv(k)=v(k)Tv(k)=v(-k). In fact, this would lead to $Tv(0)=v(0)$ and consequently to $v(0)=T^{2}v(0)=v(0)$ which is absurd. This simpler symmetry relation can instead be imposed when $T^{2}=\mathds{1}$, as discussed in [cornean2016construction] and [fiorenza2016z].
The existence of a symmetric and periodic frame is not guaranteed in general for a time-reversal symmetric projection-valued map $P$, and if $d=2$ the obstruction to constructing one, as explained in Section 4, is measured by the $\mathbb{Z}_{2}$ invariant associated with $P$. In Theorem 5.1 we will show that, instead, it is always possible to construct a pseudo-periodic symmetric frame in the sense of the following definition.
Definition 1.6 (Pseudo–periodic symmetric frame).
Let $P$ be a rank-$2n$ time-reversal symmetric projection-valued map in the sense of Definition 1.4. A frame $\{v_{j}\}_{1\leq j\leq 2n}$ is called pseudo-periodic if the vectors define continuous maps from $\pi,\pi]\times\mathbb{T}^{1}$ to $\mathcal{H}$ that satisfy
  1. (1)
    periodicity on 2n22n-2 elements, i.e.
    $v_{j}(\pi,k_{2})=v_{j}(-\pi,k_{2})\quad\forall k_{2}\in\mathbb{T}^{1},\qquad\forall j\notin\{1,n+1\},$
  2. (2)
    and periodicity up to a phase on $2$ elements, i.e. for some $h\in\mathbb{Z}$
    $v_{1}(\pi,k_{2})=e^{ihk_{2}}v_{1}(-\pi,k_{2})\quad\text{and}\quad v_{n+1}(\pi,k_{2})=e^{-ihk_{2}}v_{n+1}(-\pi,k_{2})\quad\forall k_{2}\in\mathbb{T}^{1}.$
A pseudo-periodic frame is called symmetric if the vectors further satisfy
  1. (3)
    time-reversal symmetry, i.e.
    $Tv_{i}(t,k_{2})=v_{n+i}(-t,-k_{2}),\quad\forall t\in[-\pi,\pi],\ k_{2}\in\mathbb{T}^{1}.$
In Theorem 5.1, the existence of a pseudo-periodic symmetric frame will be related to the existence of a decomposition of the projection-valued map $P$ that preserves the symmetry in an appropriate sense, namely
Definition 1.7 (Symmetric splitting).
@@
Let $P$ be a rank-$2n$ time-reversal symmetric projection-valued map in the sense of Definition 1.4. A symmetric splitting of $P$ is a pair of continuous projection-valued maps $P^{\pm}:\mathbb{T}^{d}\to\operatorname{Proj}_{n}(\mathcal{H})$ such that
$P^{+}(k)P^{-}(k)\equiv 0,\quad P(k)=P^{+}(k)+P^{-}(k)\quad\mbox{and}\quad TP^{+}(k)=P^{-}(-k)T\quad\forall k\in\mathbb{T}^{d}.$
The usual equivalence relations for projection-valued maps, which typically arise in the study of $K$-theory [rordam2000introduction], can be adapted to the time-reversal symmetric setting. It is with respect to these adapted equivalence relations that we will study the completeness of the invariant defined in Section 4.
Definition 1.8 (Equivalence relations for time-reversal symmetric projection-valued maps).
Let $P_{1},P_{2}$ be two time-reversal symmetric projection-valued maps in the sense of Definition 1.4. We say that $P_{1}$ and $P_{2}$ are:
  1. (1)
    Murray–von Neumann equivalent if there exists a continuous map $V:\mathbb{T}^{d}\to\mathcal{B}(\mathcal{H})$ with values in bounded operators such that
    $TV(k)=V(-k)T,\quad V(k)^{*}V(k)=P_{0}(k),\quad V(k)V(k)^{*}=P_{1}(k),\quad\forall k\in\mathbb{T}^{d}$
    (namely, $V(k)$ is a partial isometry between the ranges of $P_{0}(k)$ and $P_{1}(k)$ compatible with time-reversal symmetry);
  2. (2)
    unitarily equivalent if there exists a continuous unitary-valued map $U:\mathbb{T}^{d}\to\mathcal{U}(\mathcal{H})$ such that
    $TU(k)=U(-k)T,\quad P_{1}(k)=U(k)P_{0}(k)U(k)^{-1},\quad\forall k\in\mathbb{T}^{d};$
  3. (3)
    homotopic if there exists a continuous projection-valued map $P:[0,1]\times\mathbb{T}^{d}\to\operatorname{Proj}_{m}(\mathcal{H})$ such that
    $P(0,k)=P_{0}(k),\quad P(1,k)=P_{1}(k),\quad TP(t,k)=P(t,-k)T,\quad\forall t\in[0,1],\ k\in\mathbb{T}^{d}.$
Remark 1.9 (Role of Murray–von Neumann, unitary and homotopy equivalences).
As mentioned above, the three notions introduced in the previous Definition are natural equivalence relations that can be defined for projections in any $C^{*}$-algebra. In the present setting, the relevant $C^{*}$-algebra is the one of maps $A(k)$ from the torus $\mathbb{T}^{2}$ to $\mathcal{B}(\mathcal{H})$ satisfying a symmetry constraint of the form
$TA(-k)=A(k)T,\quad\forall k\in\mathbb{T}^{2}\,.$
Under suitable assumptions (for instance, when $\mathcal{H}$ is infinite-dimensional) these equivalence notions coincide. In general, however, unitary equivalence corresponds to Murray–von Neumann equivalence of both the projections and their orthogonal complements, while homotopy equivalence corresponds to unitary equivalence implemented by unitaries homotopic to the identity – unless the unitary group of the $C^{*}$-algebra is connected (which is indirectly proven in Lemma 5.5), homotopy is strictly finer than the previous two. For a detailed discussion of the hierarchy among these three equivalence relations, valid for any $C^{*}$-algebra, we refer the reader to [rordam2000introduction, Section 2.2].
@@
From a differential-geometric viewpoint, Murray–von Neumann equivalences correspond precisely to isomorphisms between the vector bundles associated with the projection-valued maps: physically, they would correspond to changes of representation on the Hilbert space of the quantum system which affect only states in the occupied energy levels, below the spectral gap of the Hamiltonian. Instead, unitary equivalences are physically associated, for example, with the different choices of dimerization (that is, of the Wigner–Seitz cell) in tight-binding models, or with the selection of a unitary-valued map that removes the $\tau$-covariance in continuum models. As for homotopies, they correspond to continuous deformations of physical parameters in the original Hamiltonian that do not close the spectral gaps. Under such deformations, the topological properties of $P$ are preserved. Since the physical interpretations of the last two mathematical notions differ, it is natural to treat them separately. Moreover, when the two conditions coincide mathematically, they establish an equivalence between distinct physical aspects of the system.
Remark 1.10 (Continuous vs. analytic frames).
Since the family of projections $\{P(k)\}_{k\in\mathbb{R}^{d}}$ arising in physical models is usually analytic in $k$, one may ask whether (real-)analytic frames can be constructed. This enhanced regularity on the frames would correspond, in position space, i.e. via the inverse Bloch–Floquet(–Zak) transform, to exponential localization of the corresponding Wannier functions, as mentioned in the Introduction. Following the argument of [cornean2016construction, Lemma 2.3], one can show that whenever a continuous frame exists, it can be modified to a real-analytic one while preserving the time-reversal symmetry.
2. Continuous parallel transport
From here on out, we focus on $2$-dimensional systems, and set $d=2$ once and for all.